OpenStudy (anonymous):

one more: 2^4 x 4^6 x 8^8 :)

OpenStudy (anonymous):

lol 2^4 x (2^2)^6 x(2^3)^8 2^4x 2^12 x 2^24 2^40= 2^(2x20) = 4^20

OpenStudy (anonymous):

really is that the answer? In my maths work the multiple choice answers don't have that answer D:

OpenStudy (anonymous):

1.099511627776 × 10^12

OpenStudy (anonymous):

If I am reading that correctly...I'm assuming that you are using 'X' to represent multiplication and not as a variable.

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

um, answer could be varies...what are your choices?

OpenStudy (anonymous):

2^40 = 4^20 = 16^10=....

OpenStudy (anonymous):

a. 2^18 b. 2^30 c. 4^18 d. 4^24 e. 16^10

OpenStudy (anonymous):

here is that, look like we have the answer: 2^20 = 4^20 = 4^(2x10) = 16^10

OpenStudy (anonymous):

:D OH! thank you for your help again!